# Target Aggregate Data Adjustment Method for Transportability Analysis Utilizing Summary-Level Data from the Target Population

### v1

@@ 5 Results
In simulation study, we have evaluated the performance of TADA under different strategies for censoring. Tables 3 to 5 summarize key simulation results for both the censoring-adjusted and censoring-ignored approaches across all scenarios under various overall censoring proportions. Specifically, each table compares the estimated bias and coverage relative to the pseudo true hazard ratio. Figure 1 provides the distribution of the overall censoring rate of each scenario under 20% and 30% average overall censoring, and the censoring rates of 40% and 50% for Figure 2. Figure 3 presents the censoring rates for the treatment and control groups across various scenarios under overall censoring rates of 20% and 30%, while Figure 4 illustrates the censoring rates in each group across scenarios under average overall censoring rates of 40% and 50%.
@@
In general, TADA with censoring adjustments demonstrates superior performance in terms of bias and coverage compared to the unadjusted approach across various censoring scenarios. As the difference in censoring rates between treatment and control groups grows (e.g., in Scenarios 3, 4, 5 and 6), the benefits of the censoring-adjusted method become more evident across overall censoring rates ranging from 20% to 50%. Compared to the true hazard ratio, the censoring-adjusted TADA consistently yields lower bias and improved coverage in most cases.
At lower overall censoring rates (20% and 30%), the advantages of the censoring-adjusted TADA are particularly prominent. For instance, under a 20% censoring rate, the censoring-adjusted hazard ratio estimates show a marked reduction in bias compared to the unadjusted estimates—such as in Scenario 1, where the adjusted bias is 0.037 versus 0.128 for the unadjusted approach, roughly one-fourth of the unadjusted bias. Significant improvements are seen in Scenarios 3 and 4, with censoring-adjusted bias estimates close to zero compared to more than twenty times the results without adjustment. Similarly, in Scenario 6, the censoring-adjusted bias decreases from 0.085 to 0.003. Additionally, the censoring-adjusted method achieves coverage exceeding 92% in most scenarios at these lower censoring rates, indicating robust performance.
As the overall censoring rate rises to 40%, the censoring-adjusted TADA continues to demonstrate lower bias than the unadjusted approach, though both methods show some decline in performance. In Scenario 3, for example, the adjusted bias is 0.139 compared to an unadjusted bias of 0.218, both exceeding 10%. However, the coverage of the censoring-adjusted estimates remains higher than that of the unadjusted estimates, achieving 81.6% versus 77.1% in Scenario 3, which underscores the robustness of the adjusted method under moderate censoring conditions.
At an overall censoring rate of 50%, the benefits of censoring adjustment, though somewhat reduced, are still present. For instance, in Scenario 9, the censoring-adjusted bias is 0.190 compared to 0.196 for the unadjusted approach. While both methods experience coverage reductions at this level, the censoring-adjusted TADA maintains a higher coverage rate, such as 80.3% versus 78.5% in Scenario 8, demonstrating its relative resilience even under more challenging censoring conditions.
As shown in Figure 3 and 4, in all scenarios, the censoring proportions between the treatment and control groups do not exhibit extreme discrepancies, meaning there is no scenario where one group had a very low censoring rate while the other had a very high rate. This balanced design across scenarios ensures that the adjustment results from TADA are broadly applicable and robust.
@@
Figure 1: The boxplots of overall censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 20% and (B) 30%. The numbers aligned with each boxplot from left to right represent the minimum, mean and maximum of the overall censoring proportion among 1000 simulation replicates.
Figure 2: The boxplots of overall censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 40% and (B) 50%. The numbers aligned with each boxplot from left to right represent the minimum, mean and maximum of the overall censoring proportion among 1000 simulation replicates.
@@
Figure 3: The boxplots of treatment (bright orange) and control (light blue) group censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 20% and (B) 30%. The numbers aligned with each boxplot from bottom to top represent the minimum (blue), mean (black) and maximum (red) of corresponding censoring proportion among 1000 simulation replicates.
Figure 4: The boxplots of treatment (bright orange) and control (light blue) group censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 40% and (B) 50%. The numbers aligned with each boxplot from bottom to top represent the minimum (blue), mean (black) and maximum (red) of corresponding censoring proportion among 1000 simulation replicates.
@@
1 0.855 (0.619, 1.435) 0.946 (0.627, 1.440) 0.818 (0.809, 0.828) 0.037 (-0.266, 0.433) 0.128 (-0.188, 0.526) 0.926 0.925
2 0.774 (0.559, 1.304) 0.862 (0.569, 1.312) 0.753 (0.743, 0.762) 0.021 (-0.251, 0.387) 0.109 (-0.189, 0.506) 0.925 0.918
3 0.587 (0.425, 1.009) 0.666 (0.436, 1.020) 0.591 (0.582, 0.601) -0.004 (-0.220, 0.263) 0.075 (-0.150, 0.367) 0.935 0.923
4 0.620 (0.448, 1.062) 0.700 (0.458, 1.072) 0.623 (0.613, 0.633) -0.003 (-0.247, 0.296) 0.077 (-0.184, 0.391) 0.943 0.932
5 0.815 (0.592, 1.374) 0.907 (0.600, 1.380) 0.788 (0.779, 0.798) 0.026 (-0.273, 0.405) 0.119 (-0.189, 0.505) 0.924 0.916
6 0.626 (0.454, 1.070) 0.709 (0.465, 1.082) 0.624 (0.614, 0.633) 0.003 (-0.234, 0.318) 0.085 (-0.167, 0.420) 0.929 0.922
7 0.749 (0.545, 1.271) 0.840 (0.555, 1.277) 0.730 (0.721, 0.740) 0.019 (-0.243, 0.381) 0.110 (-0.167, 0.489) 0.924 0.917
8 0.989 (0.706, 1.587) 1.048 (0.709, 1.585) 0.952 (0.943, 0.961) 0.037 (-0.302, 0.461) 0.097 (-0.247, 0.522) 0.938 0.940
9 1.032 (0.725, 1.589) 1.065 (0.725, 1.588) 0.990 (0.982, 0.999) 0.042 (-0.306, 0.506) 0.075 (-0.267, 0.515) 0.950 0.947
10 0.901 (0.636, 1.415) 0.946 (0.639, 1.415) 0.796 (0.787, 0.805) 0.105 (-0.196, 0.527) 0.149 (-0.153, 0.554) 0.873 0.869
Table 2: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 20%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
@@
Table 3: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 30%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
1 1.067 (0.731, 1.890) 1.174 (0.751, 1.896) 0.818 (0.809, 0.828) 0.249 (-0.176, 0.830) 0.356 (-0.067, 0.914) 0.708 0.680
2 0.962 (0.652, 1.705) 1.059 (0.673, 1.713) 0.753 (0.743, 0.762) 0.209 (-0.174, 0.728) 0.306 (-0.082, 0.844) 0.749 0.711
3 0.730 (0.487, 1.314) 0.810 (0.508, 1.331) 0.591 (0.582, 0.601) 0.139 (-0.150, 0.558) 0.218 (-0.086, 0.666) 0.816 0.771
4 0.770 (0.516, 1.385) 0.854 (0.537, 1.399) 0.623 (0.613, 0.633) 0.147 (-0.171, 0.563) 0.231 (-0.108, 0.649) 0.791 0.744
5 1.024 (0.697, 1.806) 1.124 (0.718, 1.813) 0.788 (0.779, 0.798) 0.236 (-0.174, 0.807) 0.336 (-0.083, 0.892) 0.724 0.690
6 0.804 (0.535, 1.428) 0.886 (0.557, 1.442) 0.624 (0.614, 0.633) 0.181 (-0.166, 0.642) 0.262 (-0.085, 0.714) 0.763 0.709
7 0.952 (0.644, 1.680) 1.045 (0.665, 1.690) 0.730 (0.721, 0.740) 0.222 (-0.167, 0.757) 0.315 (-0.069, 0.844) 0.737 0.692
8 1.136 (0.750, 1.895) 1.180 (0.763, 1.886) 0.952 (0.943, 0.961) 0.184 (-0.257, 0.842) 0.229 (-0.202, 0.812) 0.868 0.862
9 1.098 (0.718, 1.791) 1.124 (0.726, 1.780) 0.990 (0.982, 0.999) 0.108 (-0.316, 0.753) 0.134 (-0.270, 0.712) 0.928 0.923
10 0.993 (0.654, 1.640) 1.033 (0.665, 1.635) 0.796 (0.787, 0.805) 0.197 (-0.199, 0.767) 0.237 (-0.159, 0.776) 0.816 0.805
@@
Table 4: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 40%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
1 1.269 (0.832, 2.358) 1.386 (0.860, 2.353) 0.818 (0.809, 0.828) 0.451 (-0.093, 1.239) 0.567 (0.020, 1.325) 0.520 0.473
2 1.135 (0.738, 2.114) 1.240 (0.763, 2.112) 0.753 (0.743, 0.762) 0.382 (-0.099, 1.115) 0.487 (0.010, 1.191) 0.576 0.526
3 0.868 (0.543, 1.636) 0.945 (0.570, 1.647) 0.591 (0.582, 0.601) 0.277 (-0.117, 0.831) 0.354 (-0.034, 0.909) 0.668 0.606
4 0.911 (0.574, 1.711) 0.991 (0.600, 1.716) 0.623 (0.613, 0.633) 0.288 (-0.132, 0.932) 0.368 (-0.053, 0.961) 0.665 0.598
5 1.220 (0.795, 2.259) 1.330 (0.822, 2.253) 0.788 (0.779, 0.798) 0.432 (-0.082, 1.216) 0.542 (0.013, 1.293) 0.521 0.476
6 0.964 (0.605, 1.787) 1.044 (0.631, 1.791) 0.624 (0.614, 0.633) 0.340 (-0.106, 0.982) 0.420 (-0.000, 1.043) 0.582 0.523
7 1.135 (0.732, 2.108) 1.235 (0.760, 2.102) 0.730 (0.721, 0.740) 0.404 (-0.099, 1.143) 0.505 (0.017, 1.191) 0.530 0.481
8 1.283 (0.787, 2.221) 1.301 (0.804, 2.196) 0.952 (0.943, 0.961) 0.331 (-0.209, 1.207) 0.349 (-0.148, 1.087) 0.803 0.785
9 1.180 (0.714, 2.014) 1.186 (0.727, 1.992) 0.990 (0.982, 0.999) 0.190 (-0.301, 0.959) 0.196 (-0.243, 0.859) 0.912 0.902
10 1.083 (0.666, 1.878) 1.113 (0.683, 1.859) 0.796 (0.787, 0.805) 0.287 (-0.193, 0.995) 0.317 (-0.138, 0.931) 0.784 0.757

### v2

@@ 5 Results
We evaluated the performance of TADA under different strategies for censoring. For low to medium overall censoring proportions of 20% and 30%, Figure 1 provides the distribution of the overall censoring rate for each scenario, while Figure 2 presents the censoring rates for the treatment and control groups across various scenarios. Tables 3 and 3 summarize key simulation results for both the censoring-adjusted and censoring-ignored cases across all scenarios. Specifically, each table compares the estimated bias and coverage relative to the pseudo-true hazard ratio. The illustrations of the overall censoring rate distribution, as well as the simulation results under relatively high overall censoring proportions of 40% and 50%, are provided in the Supplementary Materials S3.
@@
In general, TADA with censoring adjustments demonstrates improved efficiency in terms of bias and coverage compared to the unadjusted approach across various censoring scenarios. Compared to the true hazard ratio, the censoring-adjusted TADA consistently yields lower bias and improved coverage in most cases. As the difference in censoring rates between treatment and control groups grows (e.g., in Scenarios 3, 4, and 6), the benefits of the censoring adjustment become increasingly evident across all tested levels of overall censoring.
At lower censoring levels, the advantages of the censoring-adjusted TADA approach are prominent. Specifically, with a 20% censoring rate, the adjusted HR estimates exhibit lower bias than those unadjusted. For example, in Scenario 1, the bias associated with the censoring-adjusted method is 0.040 (95% CI: –0.256, 0.449), which is less than half of the bias under the censoring-ignored method (0.114; 95% CI: –0.205, 0.572) and with a narrower CI. This performance gap is further amplified in scenarios with greater censoring imbalances between treatment groups. In Scenario 3, the adjusted method yields a nearly unbiased point estimate, while the unadjusted estimate incurs a bias of eight times higher in magnitude. Similarly, in Scenario 4, the bias is reduced from 0.076 to 0.022, and in Scenario 6, from 0.086 to 0.023, when censoring is properly adjusted. The censoring-adjusted strategy also improves coverage. At 20% overall censoring, nine out of ten scenarios attain coverage above 90%, compared to six out of ten falling in the unadjusted cases. The improvement is more prominent when overall censoring increases to 30%, though there is a slight decline in the absolute values of coverage.
When the overall censoring rate reaches 40% and 50%, both the censoring-adjusted and unadjusted estimators exhibit larger bias and wider confidence intervals due to severe data loss. Nevertheless, TADA with censoring adjustment continues to show substantially lower bias—often around 30%–50% less—compared to the unadjusted approach, and its coverage generally remains in a higher range. By contrast, coverage for the unadjusted estimator can drop well below 70%, reaching as low as 50% in certain cases. Full details, including scenario-specific hazard ratios, bias, and coverage, are provided in Supplementary Materials S3.
These findings confirm that although high overall censoring inevitably compromises estimation precision, accounting for censoring through proper adjustment is critical to preserving inferential validity. The censoring-adjusted TADA approach consistently outperforms the unadjusted method in both bias control and coverage maintenance, even under severe data loss conditions. Moreover, as shown in Figure 2 and S2, the censoring proportions between treatment and control never diverge to an extreme. This balanced design corresponds to many real-world scenarios where group-wise censoring differences remain moderate, ensuring TADA’s adjustment results are applicable to commonly encountered settings.
@@
Figure 1: The boxplots of overall censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 20% and (B) 30%. The numbers aligned with each boxplot from left to right represent the minimum, mean and maximum of the overall censoring proportion among 500 simulation replicates.
@@
Figure 2: The boxplots of treatment (bright orange) and control (light blue) group censoring proportion for scenarios 1 to 10 when the average overall censoring proportion among all scenarios is around (A) 20% and (B) 30%. The numbers aligned with each boxplot from bottom to top represent the minimum (blue), mean (black) and maximum (red) of corresponding censoring proportion among 500 simulation replicates.
@@
1 0.858 (0.575, 1.295) 0.932 (0.624, 1.411) 0.818 (0.811, 0.826) 0.040 (–0.256, 0.449) 0.114 (–0.205, 0.572) 0.936 0.910
2 0.797 (0.536, 1.191) 0.865 (0.580, 1.299) 0.753 (0.746, 0.760) 0.044 (–0.222, 0.365) 0.112 (–0.173, 0.486) 0.938 0.904
3 0.598 (0.400, 0.905) 0.647 (0.429, 0.996) 0.591 (0.586, 0.597) 0.007 (–0.208, 0.260) 0.056 (–0.183, 0.351) 0.934 0.932
4 0.645 (0.430, 0.964) 0.699 (0.464, 1.058) 0.623 (0.617, 0.629) 0.022 (–0.192, 0.328) 0.076 (–0.171, 0.428) 0.928 0.904
5 0.828 (0.555, 1.244) 0.902 (0.604, 1.363) 0.789 (0.782, 0.796) 0.040 (–0.267, 0.430) 0.113 (–0.214, 0.551) 0.936 0.878
6 0.647 (0.433, 0.974) 0.710 (0.473, 1.083) 0.624 (0.618, 0.630) 0.023 (–0.190, 0.319) 0.086 (–0.145, 0.432) 0.948 0.898
7 0.771 (0.516, 1.164) 0.844 (0.563, 1.281) 0.730 (0.724, 0.737) 0.041 (–0.229, 0.382) 0.113 (–0.186, 0.465) 0.944 0.896
8 1.014 (0.694, 1.535) 1.059 (0.724, 1.596) 0.952 (0.943, 0.960) 0.063 (–0.266, 0.503) 0.107 (–0.252, 0.590) 0.930 0.914
9 1.058 (0.722, 1.580) 1.079 (0.741, 1.600) 0.990 (0.981, 0.999) 0.068 (–0.294, 0.540) 0.089 (–0.285, 0.553) 0.908 0.910
10 0.923 (0.622, 1.384) 0.959 (0.654, 1.425) 0.796 (0.789, 0.803) 0.127 (–0.179, 0.502) 0.163 (–0.154, 0.551) 0.892 0.846
Table 2: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 20%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo-true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo-true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo-true hazard ratio; coverage of pseudo-true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo-true hazard ratio when adopting censoring-ignored strategy.
@@
Table 3: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 30%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo-true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo-true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo-true hazard ratio; coverage of pseudo-true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo-true hazard ratio when adopting censoring-ignored strategy.
6 Sensitivity Analysis
To fully evaluate the proposed TADA method, we perform comprehensive sensitivity analyses in several aspects, including the truncation strategies of the final weights, the sample size of the study data, and possible abnormal values in the target population, corresponding to the various elements in the process that may affect the performance of TADA. Without loss of generality, the following simulations are conducted with the relevant settings of Scenario 1 in Section 4.1 as a baseline, with overall censoring rate of 20%. Each case simulates 500 replicates and each replicate runs 200 bootstrap resamples.
6.1 Final Weights Truncation Strategy
To assess the impact of final weight truncation on the performance of TADA, we vary the quantile cutoffs applied to the final weights, ranging from 80% to 99%, alongside a no-truncation baseline. As shown in Table 7, in the censoring-ignored setting where only participation weights are applied, the bias decreases as the truncation becomes stricter, and the confidence intervals correspondingly narrow. This reflects the dominant influence of extreme weights derived from covariate imbalance, which are progressively mitigated under tighter cutoffs. In the censoring-adjusted setting where the final weights incorporate both participation and censoring components, the trend of bias exhibits a convex pattern: it is reduced at moderate cutoffs but may increase under excessively loose or tight truncation. This may arise from competing effects—retaining large IPCW values inflates variance and bias, while aggressive truncation may underweight censored individuals and induce instability. Despite this, the censoring-adjusted approach maintains lower bias and higher coverage across all settings, indicating a consistent robustness. These results suggest that a proper truncation is a useful safeguard to improve estimator reliability in practice.
6.2 Source Sample Size and Bias Source
To better understand the source and behavior of estimation bias of TADA, we increase the source sample size stepwise while holding all other aspects of the simulation design fixed. Specifically, we vary the source sample size $N_{\text{s}}\in\{300,500,800,1000\}$, while maintaining the same baseline covariate distributions, censoring mechanism, and final weight truncation threshold across all settings. According to Table 7, the bias decreases gradually from $N_{\textit{s}}=300$ to $1000$ and the confidence intervals all include zero, suggesting modest bias. Larger datasets may reduce this bias further, but we limit $N_{\textit{s}}$ to $1000$ considering the computational cost under time-varying weighting. The coverage results of censoring-adjusted cases are robust with the increase of source sample size, which supports the consistency of TADA. The close alignment between Monte Carlo standard deviations and bootstrap standard errors indicates relatively accurate variance estimation; hence, some coverage shortfalls shown in Section 5 may reflect small-sample bias or censoring effects rather than variance underestimation. Alternative resampling schemes, such as stratified or weighted bootstrap, may be more suitable than the naive bootstrap in large-sample settings with substantial censoring.
6.3 Abnormal Values in the Target Population
To evaluate the robustness of the TADA in the presence of potential abnormal values in the target population, we conduct a series of scenarios designed to mimic practical data irregularities that may compromise the reliability of sample moments. As shown in Table 4, we consider seven predefined scenarios, each introducing distinct forms of abnormality into the target population data. In the clean baseline scenario (Type 1), no modification is applied. In Types 2 and 3, we introduce binary noise by randomly selecting 1% or 5% of individuals and flipping the values of the binary covariate $X_{2}$, simulating potential misclassification or data entry errors. Types 4 and 5 simulate extreme outliers in the continuous covariate $X_{3}$ by adding several times of original standard deviations to 1% or 5% of randomly selected observations, capturing heavy-tailed or contaminated measurement scenarios. Types 6 and 7 combine both mentioned mechanisms, simultaneously introducing flips in $X_{2}$ and extreme values in $X_{3}$ at corresponding proportions.
ABV Type Proportion of Flipped in $X_{2}$ Proportion of Extreme Values in $X_{3}$ Extreme Multiplier of $X_{3}$
1
2 1%
3 5%
4 1% 3x
5 5% 3x
6 1% 1% 3x
7 5% 5% 3x
Table 4: Abnormal value pattern types in the simulated target population. “Proportion of Flipped” means the proportion of binary variable values that appear to have flip-flopped from their true values. “Proportion of Extreme Values” means the proportion of continuous variable values that are extremely skewed compared to the overall distribution. “Extreme Multiplier” means the multiplier used to construct the continuous variable outliers. The hyphen indicates the corresponding parameter setting is not in effect. ABV Abnormal value.
According to Table 7, when existing outliers in the target population, the TADA estimates show a certain degree of increase in bias and widening of the confidence intervals as the form of the outliers became more complex from a single type to the presence of a superposition of both flipped and polarized values. However, the degradation of performance is limited and the coverage and bias levels remain within acceptable limits, demonstrating overall robustness. Overall, TADA is able to control bias when faced with moderate extreme values in the target population data.
Trunc. cutoff HR (95% CI): Censor HR (95% CI): Non-censor HR (95% CI): True Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor
80% 0.886 (0.628, 1.259) 0.925 (0.657, 1.317) 0.818 (0.811, 0.826) 0.068 (–0.191, 0.439) 0.107 (–0.170, 0.492) 92.00% 88.80%
90% 0.869 (0.600, 1.266) 0.926 (0.640, 1.355) 0.818 (0.811, 0.826) 0.050 (–0.221, 0.448) 0.108 (–0.180, 0.509) 93.20% 90.80%
95% 0.858 (0.575, 1.295) 0.932 (0.624, 1.411) 0.818 (0.811, 0.826) 0.040 (–0.256, 0.449) 0.114 (–0.205, 0.572) 93.60% 91.00%
99% 0.849 (0.547, 1.353) 0.950 (0.609, 1.508) 0.818 (0.811, 0.826) 0.031 (–0.294, 0.524) 0.131 (–0.219, 0.644) 94.00% 90.20%
No Truncation 0.861 (0.545, 1.388) 0.973 (0.617, 1.546) 0.818 (0.811, 0.826) 0.043 (–0.307, 0.631) 0.155 (–0.240, 0.764) 92.60% 86.80%
@@ 6.3 Abnormal Values in the Target Population
Table 5: The simulation results of various truncation strategies for $w^{\textbf{final}}$. The columns from left to right represent truncation cutoff values of $w^{\textbf{final}}$; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo-true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo-true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo-true hazard ratio; coverage of pseudo-true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo-true hazard ratio when adopting censoring-ignored strategy. All the other settings in the simulations are identical.
$N_{\textit{s}}$ Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor Monte Carlo SD: Censor Bootstrap SE: Censor Monte Carlo SD: Non-censor Bootstrap SE: Non-censor
300 0.027 (–0.197, 0.339) 0.102 (–0.143, 0.435) 94.00% 89.60% 0.13% 0.14% 0.14% 0.15%
500 0.026 (–0.162, 0.259) 0.099 (–0.108, 0.353) 91.80% 81.80% 0.11% 0.10% 0.12% 0.11%
800 0.024 (–0.121, 0.176) 0.099 (–0.066, 0.266) 95.40% 77.40% 0.08% 0.08% 0.08% 0.09%
1000 0.021 (–0.104, 0.153) 0.096 (–0.050, 0.244) 93.60% 73.40% 0.07% 0.07% 0.08% 0.08%